Determine the OH− concentration for a solution where pH = 9.27. ___ × 10 ____M

Determine the OH concentration for a solution where pH = 9.27.

___ × 10 ____M
Answer #1

pOH = 14-pH = 14-9.27 = 4.73

[OH-] = 10^pOH = 10^-4.73

[OH-] = 0.00001862087

[OH-] = 1.86*!0^-5